OpenStudy (anonymous):

Write two new sets of parametric equations for y=3x-4

OpenStudy (sh3lsh):

1) y=t x=(t-4)/3 2)x=t y=3t-4

OpenStudy (anonymous):

oh so you used t as the value?

OpenStudy (anonymous):

i used 4t

OpenStudy (anonymous):

can you show me how to do it using 4t

OpenStudy (sh3lsh):

Sorry, was away. You can just plug in anything instead of t! Here's what I'm doing using 2t. You try 4t:

OpenStudy (faiqraees):

Hey isnt it going to be x=(t+4)/3

OpenStudy (sh3lsh):

You're right @FaiqRaees

OpenStudy (sh3lsh):

y=3x-4 1) y=2t 2t=3x-4 2t+4=3x (2t+4)/3=x 2)x=2t y=3*(2t)-4 y=6t-4

OpenStudy (sh3lsh):

Makes sense? Try using 4t!

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

y=x^4-1